# The numeraire e-variable and reverse information projection

### v3

@@ 1 Introduction
E-variables measure evidence against the null: the larger its realized value, the stronger the evidence. One can also use e-variables to make hard accept/reject decisions. In particular, Markov’s inequality implies that we can reject $H_{0}$ at level $\alpha$ if an e-variable $X$ exceeds $1/\alpha$, since ${\mathsf{P}}(X\geq 1/\alpha)\leq\alpha$ for each ${\mathsf{P}}\in{\mathcal{P}}$. This paper will not be concerned with such hard decisions, but just in the measure of evidence $X$ itself, which requires no prespecification of any level $\alpha$.
@@
Our main result establishes that under absolutely no restrictions or conditions on ${\mathcal{P}}$ or ${\mathsf{Q}}$, there exists a special ${\mathsf{Q}}$-almost surely unique and strictly positive e-variable $X^{*}$ that we call the numeraire e-variable, or just the numeraire. It is characterized by the property that ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]\leq 1$ for every other e-variable $X$. This in turn implies two other interpretable properties ${\mathbb{E}}_{\mathsf{Q}}[X^{*}/X]\geq 1$ and ${\mathbb{E}}_{\mathsf{Q}}[\log(X/X^{*})]\leq 0$, the second of which is log-optimality. In fact, $X^{*}$ is the numeraire if and only if it is log-optimal in this sense. The terminology derives from mathematical finance, where the numeraire portfolio is a central object analogous to the numeraire e-variable.
Although no assumptions on ${\mathcal{P}}$ or ${\mathsf{Q}}$ are needed in general, the theory simplifies if the alternative ${\mathsf{Q}}$ does not assign positive probability to any event that is null under every ${\mathsf{P}}\in{\mathcal{P}}$. In this case we say that ${\mathsf{Q}}$ is absolutely continuous with respect to ${\mathcal{P}}$, written ${\mathsf{Q}}\ll{\mathcal{P}}$. This natural generalization of absolute continuity for pairs of measures goes back to Halmos and Savage (1949) and is satisfied in most statistically relevant settings, while being significantly weaker than requiring a common dominating reference measure.
@@
Next, we associate to the numeraire a special sub-probability measure ${\mathsf{P}}^{*}\ll{\mathsf{Q}}$ defined via its density $d{\mathsf{P}}^{*}/d{\mathsf{Q}}=1/X^{*}$. This measure belongs to the bipolar of ${\mathcal{P}}$, denoted by ${\mathcal{P}}^{\circ\circ}$. The bipolar can be interpreted as the effective null hypothesis: it is the largest family of sub-probability measures that has the same set of e-variables as ${\mathcal{P}}$. These are the distributions against which e-variables for ${\mathcal{P}}$ do not provide any evidence.
@@
Section 2 defines the numeraire, establishes existence and uniqueness under no assumptions, characterizes the absolute continuity property ${\mathsf{Q}}\ll{\mathcal{P}}$, and establishes various other useful facts such as log-optimality of the numeraire. Section 3 introduces the RIPr and characterizes it via various optimality properties dual to the properties and log-optimality of the numeraire. Section 4 establishes a number of tools for finding the numeraire and the RIPr in practice. These tools are then applied to specific examples in Section 5. Section 6 goes beyond log-optimality by considering certain power utilities and the associated reverse Rényi projections. Section 7 outlines implications for composite alternative hypotheses and draws comparison with universal inference. Section 8 concludes.
@@
Given the measurable space $(\Omega,{\mathcal{F}})$, we let $F_{+}$, $M_{+}$, and $M_{1}$ denote the set of all $[0,\infty]$-valued measurable functions, nonnegative measures, and probability measures on $\Omega$, respectively. Given a null hypothesis ${\mathcal{P}}$, which is an arbitrary nonempty subset of $M_{1}$, we write ${\mathcal{E}}$ for the associated set of e-variables, i.e., the family of all $X\in F_{+}$ that satisfy ${\mathbb{E}}_{\mathsf{P}}[X]\leq 1$ for all ${\mathsf{P}}\in{\mathcal{P}}$. Infinite quantities play a key role in this paper. We generally rely on standard measure theoretic conventions such as $00\cdot x=$ 0 for $x\in[-\infty,\infty]$ and $x\cdot\infty=\infty$ for $x\in(0,\infty]$, as well as $\log(0)=-\infty$ and $\log(\infty)=\infty$. When evaluating ratios of $[0,\infty]$-valued random variables we often use the following convention, the last part of which is somewhat unusual but simplifies notation:
@@ 2 The numeraire e-variable
Our choice of terminology is inspired by the numeraire portfolio in the mathematical finance literature; see e.g. Long (1990); Becherer (2001); Karatzas and Kardaras (2007). In that literature, the defining property of the numeraire portfolio is that any positive wealth process divided by the numeraire becomes a supermartingale. In view of the gambling interpretation of e-variables (Shafer, 2021), this is precisely what is expressed by Definition 2.1. A subtle distinction is that we allow the numeraire to take the value infinity, whereas this is excluded in the finance literature.
@@
In the case of a simple null ${\mathcal{P}}=\{{\mathsf{P}}_{0}\}$ with ${\mathsf{Q}}$ absolutely continuous with respect to ${\mathsf{P}}_{0}$, the numeraire is just the likelihood ratio $X^{*}=d{\mathsf{Q}}/d{\mathsf{P}}_{0}$. Indeed, $X^{*}$ is ${\mathsf{Q}}$-almost surely strictly positive, it is an e-variable, and for any other e-variable $X$ we have ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]={\mathbb{E}}_{{\mathsf{P}}_{0}}[X]\leq 1$. In particular settings, such as when a reference measure exists and certain additional assumptions are satisfied, we will show that the numeraire is a likelihood ratio involving the ‘reverse information projection’ (Grünwald et al., 2024; Lardy et al., 2023). Even without these restrictions, $X^{*}$ can still be interpreted as a likelihood ratio. One of our goals is to establish this in full generality. This is the subject of Section 3.
@@
The existence of a numeraire is Lemma 2.9 below, which relies on the preceding Lemmas 2.7 and 2.8. Here we focus on the equivalence of the four conditions. Clearly (ii) implies (i). To see that (i) implies (ii), let $X^{*}$ be a ${\mathsf{Q}}$-almost surely finite numeraire and note that the numeraire property implies that any e-variable $X$ must be ${\mathsf{Q}}$-almost surely finite too. Next, we claim that (ii) is equivalent to (iii). Suppose (iii) fails. Then there is an event $A$ with ${\mathsf{Q}}(A)>0$ and ${\mathsf{P}}(A)=0$ for all ${\mathsf{P}}\in{\mathcal{P}}$, in which case $X=\infty\bm{1}_{A}$ is an e-variable that is not ${\mathsf{Q}}$-almost surely finite. Thus (ii) fails also. The converse follows by noting that every e-variable is finite ${\mathsf{P}}$-almost surely for all ${\mathsf{P}}\in{\mathcal{P}}$. Further, (iv) implies (ii) because under the former condition we have for any e-variable $X$ that ${\mathsf{Q}}(X=\infty)=\lim_{a\to\infty}{\mathsf{Q}}(X>a)=0$. It remains to prove that (ii) implies (iv). This more involved implication is the content of Lemma 2.6 below. ∎
@@ 3 Duality and reverse information projection
We continue to consider an arbitrary nonempty composite null hypothesis ${\mathcal{P}}$ and a simple alternative ${\mathsf{Q}}$ on a measurable space $(\Omega,{\mathcal{F}})$. There is a duality relationship between ${\mathcal{P}}$ and its set of e-variables. To develop this perspective, recall that $F_{+}$ and $M_{+}$ denote the set of all $[0,\infty]$-valued measurable functions and nonnegative measures on $\Omega$, respectively. The polar of ${\mathcal{P}}$ is defined by
${\mathcal{P}}^{\circ}=\{X\in F_{+}\colon{\mathbb{E}}_{\mathsf{P}}[X]\leq 1\text{ for all }{\mathsf{P}}\in{\mathcal{P}}\}.$
This is precisely the set of e-variables, i.e., ${\mathcal{P}}^{\circ}={\mathcal{E}}$. The bipolar of ${\mathcal{P}}$ is
${\mathcal{P}}^{\circ\circ}=\{{\mathsf{P}}\in M_{+}\colon{\mathbb{E}}_{\mathsf{P}}[X]\leq 1\text{ for all }X\in{\mathcal{P}}^{\circ}\}.$
The bipolar can be interpreted as the effective null hypothesis: it is the largest family of distributions whose set of e-variables is exactly ${\mathcal{E}}$. Thus it consists of those distributions against which e-variables in ${\mathcal{E}}$ do not provide any evidence. In particular, a ‘nontrivial’ e-variable $X$, in the sense that ${\mathbb{E}}_{\mathsf{Q}}[X]>1$, exists if and only if ${\mathsf{Q}}\notin{\mathcal{P}}^{\circ\circ}$. Because the constant $X=1$ is always an e-variable, elements of ${\mathcal{P}}^{\circ\circ}$ have total mass at most one, and sometimes strictly less than one. For example, if ${\mathsf{P}}$ belongs to ${\mathcal{P}}^{\circ\circ}$, then so does any ${\mathsf{P}}^{\prime}$ in $M_{+}$ that is dominated by ${\mathsf{P}}$ in the sense that ${\mathsf{P}}^{\prime}(A)\leq{\mathsf{P}}(A)$ for all $A$. In Section 4 we discuss more explicit descriptions of the bipolar; see Example 4.4, Lemma 4.5, and Corollary 4.6. A special case of the latter is Example 3.7 below, where the bipolar is computed for a simple null.
@@
The numeraire e-variable $X^{*}$ gives rise to an element ${\mathsf{P}}^{*}\in{\mathcal{P}}^{\circ\circ}$ defined by $d{\mathsf{P}}^{*}/d{\mathsf{Q}}=1/X^{*}$, understood as zero on $\{X^{*}=\infty\}$. Note that ${\mathsf{P}}^{*}\ll{\mathsf{Q}}$ by definition.
@@
Since $X^{*}$ is ${\mathsf{Q}}$-almost surely unique and strictly positive, ${\mathsf{P}}^{*}$ is well-defined. It belongs to ${\mathcal{P}}^{\circ\circ}$ because ${\mathbb{E}}_{{\mathsf{P}}^{*}}[X]={\mathbb{E}}_{\mathsf{Q}}[\bm{1}_{\{X^{*}<\infty\}}X/X^{*}]\leq{\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]\leq 1$ for all $X\in{\mathcal{E}}={\mathcal{P}}^{\circ}$ thanks to the numeraire property. ∎
Our results show that ${\mathsf{P}}^{*}$ can be interpreted as a reverse information projection, as we explain after Theorem 3.4 below. This concept first appeared in Csiszár and Tusnády (1984) and later in the PhD thesis of Li (1999, Chapter 4). It was used centrally in the context of e-variables in Grünwald et al. (2024), and was developed further in a recent preprint by Lardy et al. (2023). This earlier work assumes the existence of a reference measure, an assumption we are able to dispense with but nonetheless examine further in Section 4.
@@
We refer to the measure ${\mathsf{P}}^{*}$ in Lemma 3.1 as the reverse information projection (RIPr) of ${\mathsf{Q}}$ on ${\mathcal{P}}^{\circ\circ}$.
Our use of the RIPr terminology for ${\mathsf{P}}^{*}$ above is justified by what follows. Recall that the entropy of ${\mathsf{Q}}$ relative to ${\mathsf{P}}$ is defined by
@@
We will show that ${\mathsf{P}}^{*}$ minimizes the relative entropy $H({\mathsf{Q}}\mid{\mathsf{P}})$ over ${\mathcal{P}}^{\circ\circ}$. For this it will be convenient to use a slightly different expression for the relative entropy. For any ${\mathsf{P}}\in M_{+}$ we let ${\mathsf{P}}^{a}$ denote its absolutely continuous part with respect to ${\mathsf{Q}}$. One then has
@@
${\mathbb{E}}_{\mathsf{Q}}\left[X\frac{d{\mathsf{P}}^{a}}{d{\mathsf{Q}}}\right]\leq 1\text{ for all $X\in{\mathcal{E}}$ and ${\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$.}$ (3.2)
@@
Assume ${\mathsf{Q}}\ll{\mathcal{P}}$, let $X^{*}$ be the numeraire e-variable, and let ${\mathsf{P}}^{*}$ be an element of ${\mathcal{P}}^{\circ\circ}$ equivalent to ${\mathsf{Q}}$. The following statements are equivalent, where as above we write ${\mathsf{P}}^{a}$ for the absolutely continuous part of ${\mathsf{P}}$ with respect to ${\mathsf{Q}}$ (and hence ${\mathsf{P}}^{*}$).
@@
    ${\mathbb{E}}_{\mathsf{Q}}\left[\frac{d{\mathsf{P}}^{a}}{d{\mathsf{P}}^{*}}\right]\leq 1\text{ for all }{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$.
@@
    ${\mathbb{E}}_{\mathsf{Q}}\left[\log\frac{d{\mathsf{P}}^{a}}{d{\mathsf{P}}^{*}}\right]\leq 0\text{ for all }{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}.$
If any of these hold, one has the strong duality relation
${\mathbb{E}}_{\mathsf{Q}}[\log X^{*}]=\sup_{X\in{\mathcal{E}}}{\mathbb{E}}_{\mathsf{Q}}[\log X]=\inf_{{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}}H({\mathsf{Q}}\mid{\mathsf{P}})=H({\mathsf{Q}}\mid{\mathsf{P}}^{*}),$ (3.3)
@@
Property (ii) satisfied by the RIPr can be viewed as dual to the numeraire property and generalizes Property 2 in Li (1999, Theorem 4.3). The equivalence of (ii) and (iii) is analogous to the equivalence between the numeraire property and log-optimality of the numeraire established in Proposition 2.4. Finally, (3.3) states that strong duality holds between the problem of maximizing the expected logarithm over e-variables and the problem of minimizing relative entropy over the effective null. Note that ${\mathbb{E}}_{\mathsf{Q}}[\log X^{*}]$ is well-defined and nonnegative, possibly infinite, thanks to the numeraire property and the fact that $\log X^{*}\geq-1/X^{*}+1$.
The minimum relative entropy property of ${\mathsf{P}}^{*}$ justifies calling it the RIPr of ${\mathsf{Q}}$ on ${\mathcal{P}}^{\circ\circ}$ as in Definition 3.2. In the terminology of Lardy et al. (2023), property (ii) implies that the description gain of switching from ${\mathsf{P}}^{*}$ to any other element of ${\mathcal{P}}^{\circ\circ}$ is nonpositive, so that ${\mathsf{P}}^{*}$ is their universal reverse information projection of ${\mathsf{Q}}$ on ${\mathcal{P}}^{\circ\circ}$. Existing literature on the (universal) RIPr assumes the existence of a reference measure along with much stronger versions of the condition that ${\mathsf{Q}}\ll{\mathcal{P}}$. As we show in Section 4, in such settings our ${\mathsf{P}}^{*}$ coincides with existing notions of the (universal) RIPr of ${\mathsf{Q}}$ on ${\mathcal{P}}$; see Theorem 4.7 and its corollaries. However, our measure ${\mathsf{P}}^{*}$ exists in complete generality and satisfies properties normally associated with the RIPr; see Theorem 3.8 below for the fully general case. This leads us to Definition 3.2, which we argue is the natural definition of the RIPr in the absence of any assumptions.

### v4

@@ 1 Introduction
E-variables measure evidence against the null: the larger its realized value, the stronger the evidence. One can also use e-variables to make hard accept/reject decisions. In particular, Markov’s inequality implies that we can reject $H_{0}$ at level $\alpha$ if an e-variable $X$ exceeds $1/\alpha$, since ${\mathsf{P}}(X\geq 1/\alpha)\leq\alpha$ for each ${\mathsf{P}}\in{\mathcal{P}}$. This paper will not be concerned with such hard decisions, but just in the measure of evidence $X$ itself, which requires no pre-specification of any level $\alpha$.
@@
Our main result establishes that under absolutely no restrictions or conditions on ${\mathcal{P}}$ or ${\mathsf{Q}}$, there exists a special ${\mathsf{Q}}$-almost surely unique and strictly positive e-variable $X^{*}$ that we call the numeraire e-variable, or just the numeraire. It is characterized by the property that ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]\leq 1$ for every other e-variable $X$. This in turn implies two other interpretable properties ${\mathbb{E}}_{\mathsf{Q}}[X^{*}/X]\geq 1$ and ${\mathbb{E}}_{\mathsf{Q}}[\log(X/X^{*})]\leq 0$, the second of which is log-optimality. In fact, $X^{*}$ is the numeraire if and only if it is log-optimal in this sense. The terminology derives from mathematical finance, where the numeraire portfolio is a central object analogous to the numeraire e-variable; see e.g. Long (1990); Becherer (2001); Platen (2006); Karatzas and Kardaras (2007, 2021). In the standard Neyman-Pearson framework with a composite null and alternative, already Cvitanić and Karatzas (2001) observed how techniques from financial mathematics can be transferred to obtain optimal statistical tests.
Although no assumptions on ${\mathcal{P}}$ or ${\mathsf{Q}}$ are needed in general, the theory simplifies if the alternative ${\mathsf{Q}}$ does not assign positive probability to any event that is null under every ${\mathsf{P}}\in{\mathcal{P}}$. In this case we say that ${\mathsf{Q}}$ is absolutely continuous with respect to ${\mathcal{P}}$, written ${\mathsf{Q}}\ll{\mathcal{P}}$. This natural generalization of absolute continuity for pairs of measures goes back to Halmos and Savage (1949) and is satisfied in most statistically relevant settings, while being significantly weaker than requiring a common dominating reference measure. Huber and Strassen (1973) provide early insights on statistical testing in the absence of a reference measure.
@@
Next, we associate to the numeraire a special sub-probability measure ${\mathsf{P}}^{*}\ll{\mathsf{Q}}$ defined via its density $d{\mathsf{P}}^{*}/d{\mathsf{Q}}=1/X^{*}$. This measure belongs to the effective null hypothesis associated with ${\mathcal{P}}$, denoted by ${\mathcal{P}_{\textnormal{eff}}}$. The effective null is the largest family of sub-probability measures that has the same set of e-variables as ${\mathcal{P}}$. These are the distributions against which e-variables for ${\mathcal{P}}$ do not provide any evidence.
@@
Section 2 defines the numeraire, establishes existence and uniqueness under no assumptions, characterizes the absolute continuity property ${\mathsf{Q}}\ll{\mathcal{P}}$, and establishes various other useful facts such as log-optimality of the numeraire. Section 3 introduces the RIPr and characterizes it via various optimality properties dual to the properties and log-optimality of the numeraire. Section 4 establishes a number of tools for finding the numeraire and the RIPr in practice. These tools are then applied to specific examples in Section 5. Section 6 goes beyond log-optimality by considering certain power utilities and the associated reverse Rényi projections. Section 7 shows how our results imply that the method of universal inference is inadmissible. Section 8 concludes. Appendix A contains a generalized Lebesgue decomposition that we rely on in certain proofs, and Appendix B gives an overview of the relation between absolute continuity properties and finiteness of the numeraire and relative entropies.
@@
Given the measurable space $(\Omega,{\mathcal{F}})$, we let $F_{+}$, $M_{+}$, and $M_{1}$ denote the set of all $[0,\infty]$-valued measurable functions, nonnegative measures, and probability measures on ${\mathcal{F}}$, respectively. Given a null hypothesis ${\mathcal{P}}$, which is an arbitrary nonempty subset of $M_{1}$, we write ${\mathcal{E}}$ for the associated set of e-variables, i.e., the family of all $X\in F_{+}$ that satisfy ${\mathbb{E}}_{\mathsf{P}}[X]\leq 1$ for all ${\mathsf{P}}\in{\mathcal{P}}$. Infinite quantities play a key role in this paper. We generally rely on standard measure theoretic conventions such as $00\cdot x=$ 0 for $x\in[-\infty,\infty]$ and $x\cdot\infty=\infty$ for $x\in(0,\infty]$, as well as $\log(0)=-\infty$ and $\log(\infty)=\infty$. When evaluating ratios of $[0,\infty]$-valued random variables we often use the following convention, the last part of which is somewhat unusual but simplifies notation:
@@ 2 The numeraire e-variable
Our choice of terminology is inspired by the numeraire portfolio in the mathematical finance literature; see e.g. Long (1990); Becherer (2001); Platen (2006); Karatzas and Kardaras (2007, 2021). In that literature, the defining property of the numeraire portfolio is that any positive wealth process divided by the numeraire becomes a supermartingale. In view of the gambling interpretation of e-variables (Shafer, 2021), this is precisely what is expressed by Definition 2.1. A subtle distinction is that we allow the numeraire to take the value infinity, whereas this is excluded in the finance literature.
@@
In the case of a simple null ${\mathcal{P}}=\{{\mathsf{P}}_{0}\}$ with ${\mathsf{Q}}$ absolutely continuous with respect to ${\mathsf{P}}_{0}$, the numeraire is just the likelihood ratio $X^{*}=d{\mathsf{Q}}/d{\mathsf{P}}_{0}$. Indeed, $X^{*}$ is ${\mathsf{Q}}$-almost surely strictly positive, it is an e-variable, and for any other e-variable $X$ we have ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]={\mathbb{E}}_{{\mathsf{P}}_{0}}[X]\leq 1$. In particular settings, such as when a reference measure exists and certain additional assumptions are satisfied, we will show that the numeraire is a likelihood ratio involving the ‘reverse information projection’ (Grünwald et al., 2024; Lardy et al., 2024). Even without these restrictions, $X^{*}$ can still be interpreted as a likelihood ratio. One of our goals is to establish this in full generality. This is the subject of Section 3.
@@
Remark 2.5.
In some situations, for instance if $\log X^{*}$ has infinite ${\mathsf{Q}}$-expectation, it may be of interest measure evidence relative to a benchmark. By a benchmark we simply mean any ${\mathsf{Q}}$-almost surely strictly positive random variable $X_{\text{b}}$ such that $\log(X^{*}/X_{\text{b}})$ has finite ${\mathsf{Q}}$-expectation. Note that the numeraire itself is a possible benchmark. For any $X\in{\mathcal{E}}$, the log-optimality property (2.1) of $X^{*}$ yields
${\mathbb{E}}_{\mathsf{Q}}\left[\log\frac{X}{X_{\text{b}}}\right]={\mathbb{E}}_{\mathsf{Q}}\left[\log\frac{X}{X^{*}}\right]+{\mathbb{E}}_{\mathsf{Q}}\left[\log\frac{X^{*}}{X_{\text{b}}}\right]\leq{\mathbb{E}}_{\mathsf{Q}}\left[\log\frac{X^{*}}{X_{\text{b}}}\right].$
Thus the numeraire maximizes the ‘benchmarked’ quantity ${\mathbb{E}}_{\mathsf{Q}}[\log(X/X_{\text{b}})]$ over ${\mathcal{E}}$, and the optimal value is finite. This is true regardless of which benchmark is used, so the numeraire is benchmark-invariant. We thank an anonymous referee for pointing this out.
@@
The existence of a numeraire is Lemma 2.10 below, which relies on the preceding Lemmas 2.8 and 2.9. Here we focus on the equivalence of the four conditions. Clearly (ii) implies (i). To see that (i) implies (ii), let $X^{*}$ be a ${\mathsf{Q}}$-almost surely finite numeraire and note that the numeraire property implies that any e-variable $X$ must be ${\mathsf{Q}}$-almost surely finite too. Next, we claim that (ii) is equivalent to (iii). Suppose (iii) fails. Then there is an event $A$ with ${\mathsf{Q}}(A)>0$ and ${\mathsf{P}}(A)=0$ for all ${\mathsf{P}}\in{\mathcal{P}}$, in which case $X=\infty\bm{1}_{A}$ is an e-variable that is not ${\mathsf{Q}}$-almost surely finite. Thus (ii) fails also. The converse follows by noting that every e-variable is finite ${\mathsf{P}}$-almost surely for all ${\mathsf{P}}\in{\mathcal{P}}$. Further, (iv) implies (ii) because under the former condition we have for any e-variable $X$ that ${\mathsf{Q}}(X=\infty)=\lim_{a\to\infty}{\mathsf{Q}}(X>a)=0$. It remains to prove that (ii) implies (iv). This more involved implication is the content of Lemma 2.7 below. ∎
@@ 3 Duality and reverse information projection
We continue to consider an arbitrary nonempty composite null hypothesis ${\mathcal{P}}$ and a simple alternative ${\mathsf{Q}}$ on a measurable space $(\Omega,{\mathcal{F}})$. There is a duality relationship between ${\mathcal{P}}$ and its set of e-variables ${\mathcal{E}}$. To develop this perspective we make the following crucial definition, where we recall that $M_{+}$ denotes the set of all nonnegative measures on $\Omega$.
Definition 3.1.
The effective null hypothesis is the set
${\mathcal{P}_{\textnormal{eff}}}=\{{\mathsf{P}}\in M_{+}\colon{\mathbb{E}}_{\mathsf{P}}[X]\leq 1\text{ for all }X\in{\mathcal{E}}\}.$
The effective null hypothesis is the largest family of distributions whose set of e-variables is exactly ${\mathcal{E}}$. Thus it consists of those distributions against which e-variables in ${\mathcal{E}}$ do not provide any evidence. In particular, a ‘nontrivial’ e-variable $X$, in the sense that ${\mathbb{E}}_{\mathsf{Q}}[X]>1$, exists if and only if ${\mathsf{Q}}\notin{\mathcal{P}_{\textnormal{eff}}}$. Because the constant $X=1$ is always an e-variable, elements of ${\mathcal{P}_{\textnormal{eff}}}$ have total mass at most one, and sometimes strictly less than one. For example, if ${\mathsf{P}}$ belongs to ${\mathcal{P}_{\textnormal{eff}}}$, then so does any ${\mathsf{P}}^{\prime}$ in $M_{+}$ that is set-wise dominated by ${\mathsf{P}}$, meaning that ${\mathsf{P}}^{\prime}(A)\leq{\mathsf{P}}(A)$ for all $A$. In Section 4 we discuss more explicit descriptions of the effective null; see Example 4.4, Lemma 4.5, and Corollary 4.6. See also Example 3.12 below.
Remark 3.2.
There is a convex analytic interpretation of ${\mathcal{E}}$ and ${\mathcal{P}_{\textnormal{eff}}}$ as the so-called polar and bipolar of ${\mathcal{P}}$, usually denoted by ${\mathcal{P}}^{\circ}$ and ${\mathcal{P}}^{\circ\circ}$. [^4] This terminology is slightly nonstandard in that $F_{+}$, which appears in the definition of ${\mathcal{E}}$ (see the notation paragraph in the introduction), is not a subset of a locally convex topological vector space, whereas the standard theory of dual pairs works with such spaces. There are however exceptions, for example the bipolar theorems of Brannath and Schachermayer (1999) and Bartl and Kupper (2019). For some of our results we make use of this connection by invoking a suitable bipolar theorem; see the proof of Lemma 4.5.
The following observation is simple, yet key to the theory developed in this paper.
@@
The numeraire e-variable $X^{*}$ gives rise to an element ${\mathsf{P}}^{*}\in{\mathcal{P}_{\textnormal{eff}}}$ defined by $d{\mathsf{P}}^{*}/d{\mathsf{Q}}=1/X^{*}$, understood as zero on $\{X^{*}=\infty\}$. Note that ${\mathsf{P}}^{*}\ll{\mathsf{Q}}$ by definition.
@@
Since $X^{*}$ is ${\mathsf{Q}}$-almost surely unique and strictly positive, ${\mathsf{P}}^{*}$ is well-defined. It belongs to ${\mathcal{P}_{\textnormal{eff}}}$ because ${\mathbb{E}}_{{\mathsf{P}}^{*}}[X]={\mathbb{E}}_{\mathsf{Q}}[\bm{1}_{\{X^{*}<\infty\}}X/X^{*}]\leq{\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]\leq 1$ for all $X\in{\mathcal{E}}={\mathcal{P}}^{\circ}$ thanks to the numeraire property. ∎
Our results show that ${\mathsf{P}}^{*}$ can be interpreted as a reverse information projection, as we explain after Theorem 3.6 below. This concept first appeared in Csiszár and Tusnády (1984) and later in the PhD thesis of Li (1999, Chapter 4) as well as in Csiszar and Matus (2003). It was used centrally in the context of e-variables in Grünwald et al. (2024), and was developed further in a recent preprint by Lardy et al. (2024). This earlier work assumes the existence of a reference measure, an assumption we are able to dispense with but nonetheless examine further in Section 4.
@@
We refer to the measure ${\mathsf{P}}^{*}$ in Lemma 3.3 as the reverse information projection (RIPr) of ${\mathsf{Q}}$ on ${\mathcal{P}_{\textnormal{eff}}}$.
Our use of the RIPr terminology for ${\mathsf{P}}^{*}$ is justified by what follows. Recall that the entropy of ${\mathsf{Q}}$ relative to ${\mathsf{P}}$ is defined by
@@
We will show that ${\mathsf{P}}^{*}$ minimizes the relative entropy $H({\mathsf{Q}}\mid{\mathsf{P}})$ over ${\mathcal{P}_{\textnormal{eff}}}$. For this it will be convenient to use a slightly different expression for the relative entropy. For any ${\mathsf{P}}\in M_{+}$ we let ${\mathsf{P}}^{a}$ denote its absolutely continuous part with respect to ${\mathsf{Q}}$. One then has
@@
${\mathbb{E}}_{\mathsf{Q}}\left[X\frac{d{\mathsf{P}}^{a}}{d{\mathsf{Q}}}\right]\leq 1\text{ for all $X\in{\mathcal{E}}$ and ${\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}$.}$ (3.2)
@@
Assume ${\mathsf{Q}}\ll{\mathcal{P}}$ and let ${\mathsf{P}}^{*}$ be an element of ${\mathcal{P}_{\textnormal{eff}}}$ equivalent to ${\mathsf{Q}}$. The following statements are equivalent, where as above we write ${\mathsf{P}}^{a}$ for the absolutely continuous part of ${\mathsf{P}}$ with respect to ${\mathsf{Q}}$ (and hence ${\mathsf{P}}^{*}$).
@@
    ${\mathbb{E}}_{\mathsf{Q}}\left[\frac{d{\mathsf{P}}^{a}}{d{\mathsf{P}}^{*}}\right]\leq 1\text{ for all }{\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}$.
@@
    ${\mathbb{E}}_{\mathsf{Q}}\left[\log\frac{d{\mathsf{P}}^{a}}{d{\mathsf{P}}^{*}}\right]\leq 0\text{ for all }{\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}.$
Letting $X^{*}$ be the numeraire and ${\mathsf{P}}^{*}$ the RIPr, one has the strong duality relation
${\mathbb{E}}_{\mathsf{Q}}[\log X^{*}]=\sup_{X\in{\mathcal{E}}}{\mathbb{E}}_{\mathsf{Q}}[\log X]=\inf_{{\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}}H({\mathsf{Q}}\mid{\mathsf{P}})=H({\mathsf{Q}}\mid{\mathsf{P}}^{*}),$ (3.3)
@@
Property (ii) satisfied by the RIPr can be viewed as dual to the numeraire property and generalizes Property 2 in Li (1999, Theorem 4.3). The equivalence of (ii) and (iii) is analogous to the equivalence between the numeraire property and log-optimality of the numeraire established in Proposition 2.4; we elaborate on this analogy in Remark 3.7 below. Finally, (3.3) states that strong duality holds between the problem of maximizing the expected logarithm over e-variables and the problem of minimizing relative entropy over the effective null. Note that ${\mathbb{E}}_{\mathsf{Q}}[\log X^{*}]$ is well-defined and nonnegative, possibly infinite, thanks to the numeraire property and the fact that $\log X^{*}\geq-1/X^{*}+1$. In the spirit of Remark 2.5, we discuss a ‘benchmarked’ version of (3.3) in Remark 3.8 below.
The minimum relative entropy property of ${\mathsf{P}}^{*}$ justifies calling it the RIPr of ${\mathsf{Q}}$ on ${\mathcal{P}_{\textnormal{eff}}}$ as in Definition 3.4. In the terminology of Lardy et al. (2024), property (iii) states that the description gain of switching from ${\mathsf{P}}^{*}$ to any other element of ${\mathcal{P}_{\textnormal{eff}}}$ is nonpositive, and this provides another justification for calling ${\mathsf{P}}^{*}$ the RIPr. We elaborate on the connection to their work in Remark 3.9 below. Existing literature on the RIPr assumes the existence of a reference measure along with much stronger versions of the condition that ${\mathsf{Q}}\ll{\mathcal{P}}$. As we show in Section 4, in such settings our ${\mathsf{P}}^{*}$ coincides with existing notions of the RIPr of ${\mathsf{Q}}$ on ${\mathcal{P}}$; see Theorem 4.7 and its corollaries. However, our measure ${\mathsf{P}}^{*}$ exists in complete generality and satisfies properties normally associated with the RIPr; see Theorem 3.13 below for the fully general case. This leads us to Definition 3.4, which we argue is the natural definition of the RIPr in the absence of any assumptions.